Could you please give detailed, step by step value of solution and also explain what is incremental value of acquisition? Thank you very much!!!!

Moore Industries has agreed to be acquired by Scott Enterprises for 800 shares of Scott Enterprises stock. Scott Enterprises currently has 7,500 shares of stock outstanding at a price of $28 a share. Moore Industries has 1,800 shares outstanding at a price of $12 a share. The incremental value of the acquisition is $1,100. What is the value per share of Scott Enterprises stock after the acquisition?

a) 27.52

b) 27.96

c) 28.04

d) 28.47

e) 31.03
Answer #1

Solution

Scott enterprises will pay to = 800*28= $22,400

Value of shares of Scott enterprises = 7500*28= $2,10,000

Value of Moore industries = 1800*12= $21,600

Incremental value of acquisition = $1100

Value per share = (210000+21600+1100-22400)/7500

Value per share = 2,10,300/7500= $28.04

Option C is correct I.e. $28.04